OpenStudy (anonymous):

A circuit maintains a constant resistance. If the current in the circuit is doubled, what is the effect on the power dissipated by the circuit?

OpenStudy (kropot72):

The power is proportional to the square of the current. So if the current was tripled, the power would increase to nine times its original value. Can you now find the answer?

OpenStudy (unklerhaukus):

Ohm's Law \[V=IR\] Electrical Power \[P=IV=I(IR)=I^2R\]

OpenStudy (unklerhaukus):

\[P_0=I^2_0R\] \[P_1=I^2_1R=(2I_0)^2R=\{\text{simplify, and express in-terms of \(P_0\)}\} %=4I_0^2R=4P_0 \]